# DiaHalu: A Dialogue-level Hallucination Evaluation Benchmark for Large Language Models

### v1

@@ Abstract
Since large language models (LLMs) achieve significant success in recent years, the hallucination issue remains a challenge, numerous benchmarks are proposed to detect the hallucination. Nevertheless, some of these benchmarks are not naturally generated by LLMs but are intentionally induced. Also, many merely focus on the factuality hallucination while ignoring the faithfulness hallucination. Additionally, although dialogue pattern is more widely utilized in the era of LLMs, current benchmarks only concentrate on sentence-level and passage-level hallucination. In this study, we propose DiaHalu, the first dialogue-level hallucination evaluation benchmark to our knowledge. Initially, we integrate the collected topics into system prompts and facilitate a dialogue between two ChatGPT3.5. Subsequently, we manually modify the contents that do not adhere to human language conventions and then have LLMs re-generate, simulating authentic human-machine interaction scenarios. Finally, professional scholars annotate all the samples in the dataset. DiaHalu covers four common multi-turn dialogue domains and five hallucination subtypes, extended from factuality and faithfulness hallucination. Experiments through some well-known LLMs and detection methods on the dataset show that DiaHalu is a challenging benchmark, holding significant value for further research.
DiaHalu: A Dialogue-level Hallucination Evaluation Benchmark for Large Language Models
@@ 1 Introduction
Figure 1: Our benchmark not only includes factuality hallucination but also incorporates faithfulness hallucination at the dialogue level, although most benchmarks overlook the latter.
@@
Large language models (LLMs) (Zhao et al., 2023a), addressing many downstream tasks through natural language generation (NLG) technique, achieve significant success across diverse fields of natural language processing (NLP) (Li et al., 2024; Pan et al., 2024; Wei et al., 2023; Fan et al., 2023). With a considerable volume of parameters and sophisticated training methodologies, LLMs significantly propelled advancements in artificial intelligence (Zhao et al., 2023a).
@@
Despite many advantages of large language models, the issue of hallucination remains a primary concern (Ji et al., 2023; Zhang et al., 2023). Hallucination predominantly centers on the inclination of LLMs to generate nonsensical or untruthful contents for specific sources (Wang et al., 2023). Therefore, the occurrence of hallucinations poses risks to the application of large language models in various real-world scenarios (Omiye et al., 2023; Kasneci et al., 2023; Wu et al., 2023a).
@@
Given the aforementioned risks, hallucination detection emerges as a highly crucial task. In recent years, researchers propose numerous benchmarks for hallucination detection task (Li et al., 2023; Guan et al., 2023; Manakul et al., 2023; Yang et al., 2023). Nevertheless, several problems persist in these benchmarks. (1) Not naturally generated. One pitfall of existing benchmarks for detecting LLMs’ hallucination is that the hallucinated contents are typically induced via manually designed trigger prompts (Li et al., 2023), while not naturally generated by LLMs as in daily usage (Liu et al., 2022). (2) Merely focusing on factuality hallucination. Most benchmarks merely focus on detecting factuality hallucination (Guan et al., 2023), with few datasets that can demonstrate faithfulness hallucination (Huang et al., 2023a) (Figure 1). (3) Only concentrating on sentence-level and passage-level. Researchers proposes many sentence-level (Manakul et al., 2023; Zhao et al., 2023b) and passage-level (Yang et al., 2023; Feng et al., 2023b) hallucination detection benchmarks. However, the dialogue pattern has broader and more widespread applications in LLMs. Although it is not mentioned in previous researches to our current knowledge, dialogue-level hallucination detection is equally essential.
@@
Therefore, we propose a new dialogue-level hallucination evaluation benchmark for large language models (DiaHalu). We initially determine four domains for multi-turn dialogue: knowledge-grounded, task-oriented, chit-chat and reasoning. For each domain, we undertake a three-step process to construct the dataset. (1) We collect topics for dialogues from various sources. (2) We incorporate the topics into artificially designed system prompts and input them into two ChatGPT3.5, enabling them to engage in a multi-turn dialogue. (3) Since the knowledge-grounded and task-oriented domains stand for real human-machine interaction scenarios, we align the contents of one of the conversational participants with human language. We manually modify the contents that do not conform to human language conventions and have LLMs re-generate, resulting in the final outcomes. Then, professional scholars annotate all the samples with labels, hallucination types and locations, as well as explanations. It is noteworthy that we not only consider the factuality hallucination but also further classify the faithfulness hallucination into three types: Incoherence, Irrelevance and Overreliance. We similarly introduce the reasoning hallucination for the reasoning domain. Additionally, we conduct experiments on DiaHalu by deploying existing hallucination detection methods and some commonly used LLMs. The results indicate that DiaHalu is a highly challenging dataset, and both Chain-of-Thought and retrieval methods can improve the performance of the detection. Our contributions can be listed as follows:
@@
· To our current knowledge, we propose the first dialogue-level hallucination detection benchmark named DiaHalu [^1] https://anonymous.4open.science/r/DiaHalu-A1E9/.
@@
· DiaHalu covers four multi-turn dialogue domains along with five hallucination subtypes extended from factuality and faithfulness hallucinations, which are more widely applicable in real-world scenarios.
@@
· The experimental results indicate that DiaHalu is a highly challenging task for most LLMs and existing detection methods, holding significant value for further research.
@@ 2.1 Hallucination Detection Benchmarks
Hallucination predominantly centers on the inclination of large language models to generate nonsensical or untruthful contents for specific sources (Wang et al., 2023). Therefore, detecting hallucination in LLMs is of great importance. In recent years, researchers propose numerous benchmarks for the hallucination detection task.
@@
In earlier years, hallucination detection benchmarks are primarily curated through manual methods or generated via conventional language models. FactCollect (Ribeiro et al., 2022) is an artificially generated, multi-source factual hallucination detection dataset. Muhlgay et al. (2023) collects error samples by instructing the language model based on pre-defined error types. HADES (Liu et al., 2022) constitutes a hallucination detection dataset produced by conventional language model BERT (Devlin et al., 2019). These benchmarks are not naturally generated by LLMs as in daily usage.
@@
Consequently, some benchmarks are proposed to investigate the capacity for the direct generation of large language models. Zhao et al. (2023b); Fu et al. (2023); Chen et al. (2023a); Huang et al. (2023b); Zheng et al. (2023) enable LLMs to handle Question-Answer (QA) tasks and assess the factual accuracy of their responses. Concept-7 dataset used by Luo et al. (2023), evaluates whether a language model genuinely comprehends the meaning of each concept, thereby determining the presence of hallucination. FactCHD (Chen et al., 2023c) is generated based on natural language text and knowledge graphs (KGs), deploying two major steps: Query and Response Data Simulation & Chain of Evidence and Human Filter, to detect fact-conflicting hallucination in LLMs. Mündler et al. (2023) employs a generative language model (gLM) to rewrite sentences according to the given context. New sentences compose a dataset that can evaluate whether the generated sentences exhibit knowledge-based self-contradiction hallucination. The aforementioned benchmarks mainly focus on detecting factuality hallucination (Guan et al., 2023), while ignoring the faithfulness hallucination (Huang et al., 2023a). The benchmark proposed in this paper extends to include faithfulness hallucination, that is, to evaluate the coherence and relevance of generated contents by LLMs.
@@
Researchers also raise many sentence-level (Manakul et al., 2023; Zhao et al., 2023b; Guan et al., 2023) and passage-level (Yang et al., 2023; Feng et al., 2023b; Li et al., 2023) hallucination detection benchmarks. Nevertheless, the dialogue pattern holds broader applications within LLMs. While previous research does not, to our current knowledge, propose a dialogue-level hallucination detection benchmark. So, our DiaHalu is at the dialogue level.
2.2 Dialogue System in LLMs
It is the formidable dialogue capability exhibited by ChatGPT3.5 (Wu et al., 2023b) and GPT4 (OpenAI, 2023), that stimulates a heightened interest among researchers in large language models. The ability for multi-turn dialogue is among the most crucial features of LLMs, which has broader and more widespread applications in the current development of NLP.
Dialogue system can be primarily categorized into various types, including task-oriented dialogue (Wen et al., 2017), chit-chat dialogue (Sun et al., 2021), knowledge-grounded dialogue (Ghazvininejad et al., 2018), etc. In these scenarios, language models are required to exhibit long-term memory, recognize shifts in dialogue states, encompass extensive knowledge, and engage in reasoning processes (Huang and Chang, 2023). Despite the remarkable general capability of LLMs, there are still deficiencies in handling complex dialogue scenarios (Chen et al., 2017; Deriu et al., 2021).
Due to the critical significance of the dialogue ability, this paper underscores the detection of hallucination produced by LLMs in multi-turn dialogues.
@@ 2.2 Dialogue System in LLMs
Figure 2: The demonstration of the DiaHalu benchmark, which covers four domains and five hallucination types within dialogue-level scenarios. We also provide explanations and sources in the benchmark.
@@ 3.1 Principles
The primary objective of this benchmark is to conduct hallucination detection in large language models. Hence, it is imperative to comprehend the meaning of hallucination. Hallucination predominantly centers on the inclination of LLMs to generate nonsensical or untruthful contents regarding specific sources (Wang et al., 2023). The significance of hallucination detection lies in elevating the quality of text generation, preventing misleading information and misunderstandings, supporting applications within professional domains, etc. Therefore, to enhance the universality of our benchmark, it encompasses various multi-turn dialogue scenarios and multiple types of hallucination, extending from factuality hallucination and faithfulness hallucination (Huang et al., 2023a).
@@ 3.2 Hallucination on Diverse Domains
We consider hallucination in diverse domains of multi-turn dialogue scenarios. Our benchmark covers a total of four domains, which are shown in Figure 2. Below are their specific descriptions.
@@
involves open-ended and non-goal dialogues (Sun et al., 2021). We provide two LLMs with personas and facilitate a dialogue between them. This approach allows for the evaluation of their memory capabilities, conversational coherence, and relevance to the information being discussed.
@@
centralizes on the logical reasoning and understanding capabilities of LLMs. We have the models discuss mathematical problems to achieve the answers (Kakarla et al., 2024).
3.3 Hallucination Types
We consider factuality hallucination and faithfulness hallucination (Huang et al., 2023a). Based on Chen et al. (2023b), we carry out a detailed classification of faithfulness hallucination into Incoherence, Irrelevance, and Overreliance. Meanwhile, we introduce Reasoning Error within the reasoning dialogue. The example of each type can be referenced in Figure 2 and Appendix A.4.
@@ 3.3 Hallucination Types
is that the LLM excessively relies on the correctness of context, generating responses for unanswerable content (Slobodkin et al., 2023).
@@ 4.1 The Collection of Dialogue Topics
For knowledge-grounded dialogue, we take into account world knowledge, factual knowledge, commonsense knowledge and multi-hop web knowledge. Therefore, we gather dialogue topics for TruthfulQA (Lin et al., 2022), CommonsenseQA (Talmor et al., 2019) and CWQ (Talmor and Berant, 2018) datasets. There are also topics provided by GPT4 (OpenAI, 2023) and social media (including the authors). As for task-oriented dialogue, we primarily apply the most widely used MultiWOZ (MultiWOZ 2.1) (Budzianowski et al., 2018) which has 7 real-life scenarios. To enrich the dialogue settings, we also consider the DSTC (DSTC 1.0) (Williams et al., 2013) dataset with a focus on bus routes. GPT4 and social media are harnessed to augment user behaviors and generate more dialogue occasions. We define the LLMs with personas primarily from Jandaghi et al. (2023) and facilitate an open chit-chat dialogue between them. Additionally, we make use of mathematical problems to assess the logical reasoning dialogue ability of LLMs. These problems are sourced from GSM8K (Cobbe et al., 2021) and MathQA (Amini et al., 2019), both of which involve mathematical problems and solving processes encountered by middle school students. The overall distribution of the above sources for dialogue topics is illustrated in Figure 3.
Figure 3: The distribution of the 10 sources for dialogue topics.
@@ 4.2 Dialogue Generation
Once finishing collecting the dialogue topics for each domain, we leverage ChatGPT3.5 to generate conversations in the format of self-dialogue. The complete process of dialogue generation is illustrated in Figure 4.
@@
Initially, we integrate the dialogue topics into two system prompts, which are then inputted separately into two ChatGPT3.5. These two system prompts guide the LLMs to generate N rounds of dialogue in a given domain and topic. More details of the system prompts can be found in Appendix A.1. Then, for knowledge-grounded dialogue and task-oriented dialogue, we manually examine all responses from A to determine their adherence to human language. For in both scenarios, we consider real human-machine interaction, aiming to assess the LLMs’ adaptability to genuine human behaviors. (We assume that A is the user and B is the LLM. In this setup, we ensure the accuracy of A and only annotate the contents of B.) The chit-chat dialogue and reasoning dialogue are relatively unconstrained and freely conducted, without incorporating any specific human-machine interaction settings (Section 3.2). They necessitate only their memory and comprehension capabilities regarding contextual information, thereby minimizing the need for manual intervention. Consequently, for the responses of A in knowledge-grounded and task-oriented scenarios, where the contents do not conform to human language, we manually modify and have LLMs re-generate. Eventually, we obtain the complete dataset of multi-turn dialogues.
We provide one output sample in Appendix A.2. The comparison with other benchmarks is in Appendix A.3, highlighting the distinctive features that our DiaHalu is a naturally generated dialogue-level benchmark by LLMs, with various kinds of hallucination and explanations. The statistical information of the whole benchmark is in Table 1.
@@
Avg. Response Length (Words) 13.5091

### v2

@@ Abstract
Since large language models (LLMs) achieve significant success in recent years, the hallucination issue remains a challenge, and numerous benchmarks are proposed for hallucination detection. Nevertheless, some of these benchmarks are not naturally generated by LLMs but are intentionally induced. Also, many merely focus on the factuality hallucination while ignoring the faithfulness hallucination. Additionally, although dialogue pattern is more widely utilized in the era of LLMs, current benchmarks only concentrate on sentence-level and passage-level hallucination. In this study, we propose DiaHalu, the first dedicated dialogue-level hallucination evaluation benchmark for LLMs to our knowledge. Initially, we integrate the collected topics into system prompts and facilitate a dialogue between two LLMs. Subsequently, we manually modify the contents that do not adhere to human language conventions and then have LLMs re-generate, simulating authentic human-machine interaction scenarios. Finally, professional scholars annotate all the samples in the dataset. DiaHalu covers four common multi-turn dialogue domains and five hallucination subtypes, extended from factuality and faithfulness hallucination. Experiments through some well-known LLMs and detection methods on the dataset show that DiaHalu is a challenging benchmark, holding significant value for further research [^1] https://github.com/141forever/DiaHalu.
DiaHalu: A Dialogue-level Hallucination Evaluation Benchmark
for Large Language Models
@@ 1 Introduction
Figure 1: Our benchmark not only includes factuality hallucination but also incorporates faithfulness hallucination at the dialogue level, although most benchmarks overlook the latter one.
@@
Large language models (LLMs) (Zhao et al., 2023b), addressing many downstream tasks through natural language generation (NLG) technique, achieve significant success across diverse fields of natural language processing (NLP) (Li et al., 2024a; Pan et al., 2024; Wei et al., 2023; Fan et al., 2023). With a considerable volume of parameters and sophisticated training methodologies, LLMs significantly propelled advancements in artificial intelligence (Zhao et al., 2023b).
@@
Despite many advantages of large language models, the issue of hallucination remains a primary concern (Ji et al., 2023; Zhang et al., 2023c). Hallucination predominantly centers on the inclination of LLMs to generate nonsensical or untruthful contents for specific sources (Wang et al., 2023a). Therefore, the occurrence of hallucination poses risks to the application of large language models in various real-world scenarios (Omiye et al., 2023; Wu et al., 2023b; Chen et al., 2024).
@@
Given the aforementioned risks, hallucination detection emerges as a highly crucial task. In recent years, researchers propose numerous benchmarks for hallucination detection task (Li et al., 2023b; Guan et al., 2023; Manakul et al., 2023; Yang et al., 2023). Nevertheless, several problems persist in these benchmarks. (1) Not naturally generated. One pitfall of existing benchmarks for detecting LLMs’ hallucination is that the hallucinated contents are typically induced via manually designed trigger prompts (Li et al., 2023b), while not naturally generated by LLMs as in daily usage (Liu et al., 2022). (2) Merely focusing on factuality hallucination. Most benchmarks merely focus on detecting factuality hallucination (Guan et al., 2023), with few datasets that can demonstrate faithfulness hallucination (Huang et al., 2023a) (Figure 1). (3) Only concentrating on sentence-level and passage-level. Researchers propose many sentence-level (Manakul et al., 2023; Zhao et al., 2023c) and passage-level (Yang et al., 2023; Feng et al., 2023b) hallucination detection benchmarks. However, the dialogue pattern has broader and more widespread applications in LLMs. Although it is rarely mentioned in previous researches, dialogue-level hallucination detection is equally essential.
@@
Therefore, we propose a new dialogue-level hallucination evaluation benchmark for large language models (DiaHalu). We initially determine four domains for multi-turn dialogue: knowledge-grounded, task-oriented, chit-chat and reasoning. For each domain, we undertake a three-step process to construct the dataset. (1) We collect topics for dialogue from various sources, incorporate the topics into artificially designed system prompts and input them into two LLMs, enabling them to engage in a multi-turn dialogue. (2) Since the knowledge-grounded and task-oriented domains stand for real human-machine interaction scenarios, we align the contents of one of the conversational participants with human language. We manually modify the contents that do not conform to human language conventions and have LLMs re-generate, resulting in the final responses. (3) Professional scholars annotate all the samples with labels, hallucination subtypes and locations, as well as explanations. It is noteworthy that we not only consider the factuality hallucination but also further classify the faithfulness hallucination into three types: Incoherence, Irrelevance and Overreliance. We similarly introduce the reasoning hallucination for the reasoning domain. The advantages of ours compared with previous benchmarks are listed in Table 1. Additionally, we conduct experiments on DiaHalu by deploying existing hallucination detection methods and some commonly used LLMs. The results indicate that DiaHalu is a highly challenging benchmark. Our contributions can be listed as follows:
@@
  • To our current knowledge, we propose the first dedicated dialogue-level hallucination detection benchmark for LLMs named DiaHalu.
@@
  • DiaHalu covers four multi-turn dialogue domains along with five hallucination subtypes extended from factuality and faithfulness hallucination, which are more widely applicable in real-world scenarios.
@@
  • The experimental results indicate that DiaHalu is a highly challenging benchmark for most LLMs and existing detection methods, holding significant value for further researches.
Benchmark By LLMs Faith Halu Multi Dia Explanation
FactCollect - - - -
BEGIN - - $\checkmark$ $\checkmark$
HADES - $\checkmark$ - $\checkmark$
FactCHD $\checkmark$ - - $\checkmark$
HaluEval - - $\checkmark$ $\checkmark$
WikiBio+ $\checkmark$ - - -
PHD $\checkmark$ - - -
Ours $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 1: The comparison between our DiaHalu and other benchmarks. ‘By LLMs’, ‘Faith Halu’, ‘Multi Dia’, and ‘Explanation’ mean whether it is naturally generated by LLMs, whether it provides faithfulness hallucination, whether it is at multi-turn dialogue level, and whether there are explanations respectively (Appendix A.1).
@@ 2.1 Hallucination Detection Benchmarks
In recent years, researchers propose numerous benchmarks for hallucination detection.
@@
In earlier years, hallucination detection benchmarks are primarily organized through manual methods or generated via conventional language models. FactCollect (Ribeiro et al., 2022) is an artificially generated, multi-source factual hallucination detection benchmark. Muhlgay et al. (2023) collects error samples by instructing the language model based on pre-defined error types. HADES (Liu et al., 2022) and BEGIN (Dziri et al., 2022c) constitute hallucination detection datasets by conventional language model BERT (Devlin et al., 2019) and T5 (mostly) (Raffel et al., 2020) respectively. These benchmarks are not naturally generated by LLMs as in daily usage.
@@
Consequently, some benchmarks are proposed to investigate the direct generation abilities of large language models. Zhao et al. (2023c); Fu et al. (2023); Chen et al. (2023a); Huang et al. (2023b); Zheng et al. (2023) enable LLMs to handle Question-Answer (QA) task and assess the factual accuracy of their responses. Concept-7 dataset used by Luo et al. (2023) evaluates whether a language model truly comprehends the meaning of each concept, thereby determining the presence of hallucination. FactCHD (Chen et al., 2023c) is generated based on natural language text and knowledge graphs (KGs). Mündler et al. (2023) employs a generative language model (gLM) to rewrite sentences according to the given context. New sentences compose a dataset that can evaluate whether the generated sentences exhibit knowledge-based self-contradiction hallucination. The aforementioned benchmarks mainly focus on detecting factuality hallucination (Guan et al., 2023), while ignoring the faithfulness hallucination (Huang et al., 2023a). The benchmark proposed in this paper extends to include faithfulness hallucination, that is, to evaluate the coherence and relevance of contents generated by LLMs.
@@
Researchers also raise many sentence-level (Manakul et al., 2023; Zhao et al., 2023c; Guan et al., 2023) and passage-level (Yang et al., 2023; Feng et al., 2023b; Li et al., 2023b) hallucination detection benchmarks. Nevertheless, the dialogue pattern holds broader applications within LLMs. While previous researches rarely, to our current knowledge, propose a dialogue-level hallucination detection benchmark for LLMs. So, our DiaHalu is at the dialogue level.
2.2 Hallucination Detection
Current methods for hallucination detection (Tonmoy et al., 2024) can mainly be divided into four categories. (1) Model-Based. This method involves having the language models perform a classification task to determine whether hallucinated contents are present (Zhao et al., 2023a; Maharaj et al., 2023). (2) Retrieval-Based. For the limited knowledge within the parameters of language models, we can detect hallucination by extracting or retrieving relevant knowledge from external knowledge graphs (Martino et al., 2023; Chen et al., 2023c) or web information sources (Béchard and Ayala, 2024). (3) Sample-Based. Another feasible method is to rewrite the generated contents to evaluate the consistency (Manakul et al., 2023; Zhao et al., 2023c; Zhang et al., 2023a). (4) Proxy-Based. The mainstream view of this method (Zhang et al., 2023b; Gupta et al., 2024) is that ‘the lower the probability of generating a token, the more likely a model is to produce hallucination’.
@@ 2.2 Hallucination Detection
Figure 2: The demonstration of the DiaHalu benchmark, which covers four domains and five hallucination subtypes within dialogue-level scenarios. We also provide explanations and sources in the benchmark.
@@ 3.1 Principles
The primary objective of this benchmark is to conduct hallucination detection in large language models. Hence, it is imperative to comprehend the meaning of hallucination. Hallucination predominantly centers on the inclination of LLMs to generate nonsensical or untruthful contents regarding specific sources (Wang et al., 2023a). The significance of hallucination detection lies in elevating the quality of text generation, preventing misleading information and misunderstandings, supporting applications within professional domains, etc. Therefore, to enhance the universality of our benchmark, it encompasses various multi-turn dialogue scenarios and multiple subtypes of hallucination, extending from factuality hallucination and faithfulness hallucination (Huang et al., 2023a).
@@ 3.2 Hallucination on Diverse Domains
We consider hallucination in diverse domains of multi-turn dialogue scenarios. Our benchmark covers a total of four domains, shown in Figure 2. Their specific descriptions are in Appendix A.2.
@@
is designed for users to engage in knowledge-based dialogue with LLMs (Ghazvininejad et al., 2018). The two speakers take part in a conversation about a knowledge-based issue.
@@
is in a form of human-computer interaction, intending to accomplish a user-specified task (Wen et al., 2017).
@@
involves open-ended and non-goal dialogue (Sun et al., 2021). We provide two LLMs with personas and facilitate a dialogue between them.
@@
Following previous works (Chen et al., 2023b; Li et al., 2023a; Buszydlik et al., 2024; Grover et al., 2024; Zheng et al., 2023; Huang and Chang, 2023), we also treat reasoning errors as a kind of hallucination. We have the models discuss mathematical problems to achieve the answers (Kakarla et al., 2024).
3.3 Hallucination Taxonomy
We consider both factuality and faithfulness hallucination (Huang et al., 2023a). Based on Chen et al. (2023b); Wu et al. (2023a); Dziri et al. (2022b) and early works on text coherence (Wolf and Gibson, 2004; Atwell et al., 2024), we carry out a detailed classification of faithfulness hallucination into Incoherence, Irrelevance, and Overreliance. Meanwhile, we introduce Reasoning Error within the reasoning dialogue (Chen et al., 2023b; Li et al., 2023a; Buszydlik et al., 2024; Grover et al., 2024; Zheng et al., 2023; Huang and Chang, 2023). The example of each type can be referenced in Figure 2.
@@ 3.3 Hallucination Taxonomy
is that the LLM excessively trusts in the correctness of the context, generating responses for unanswerable contents (Slobodkin et al., 2023).
@@ 4.1 The Collection of Dialogue Topics
For knowledge-grounded dialogue, we take into account world knowledge, factual knowledge, commonsense knowledge and multi-hop web knowledge. Therefore, we gather dialogue topics from TruthfulQA (Lin et al., 2022), CommonsenseQA (Talmor et al., 2019) and CWQ (Talmor and Berant, 2018) datasets. There are also topics provided by GPT4 (OpenAI, 2023) and social media (including the authors). As for task-oriented dialogue, we primarily apply the most widely used MultiWOZ (MultiWOZ 2.1) (Budzianowski et al., 2018) which covers 7 real-life scenarios. To enrich the dialogue settings, we also consider the DSTC (DSTC 1.0) (Williams et al., 2013) dataset with a focus on bus routes. GPT4 and social media are harnessed to augment user behaviors and generate more dialogue occasions. We define the LLMs with personas primarily from Jandaghi et al. (2023) and facilitate an open chit-chat dialogue between them. Additionally, we make use of mathematical problems to assess the logical reasoning dialogue abilities of LLMs. These problems are sourced from GSM8K (Cobbe et al., 2021) and MathQA (Amini et al., 2019), both of which involve mathematical problems and solving processes encountered by middle school students.
The overall distribution of the above sources for dialogue topics is illustrated in Appendix A.3.
@@ 4.2 Dialogue Generation
Once finishing collecting the dialogue topics for each domain, we leverage ChatGPT3.5 and GPT4 to generate conversations in the format of self-dialogue. The complete process of dialogue generation is illustrated in Figure 3.
Initially, we integrate the dialogue topics into two system prompts, which are then inputted separately into two LLMs (both are ChatGPT3.5 or GPT4). These two system prompts guide the LLMs to generate $N$ rounds of dialogue in a given domain and topic. More details of the system prompts can be found in Appendix A.4. Then, for knowledge-grounded dialogue and task-oriented dialogue, we manually examine all responses from A to determine their adherence to human language. For in both scenarios, we consider real human-machine interaction, aiming to assess the LLMs’ adaptability to genuine human behaviors. (We assume that A is the user and B is the LLM. In this setup, we ensure the accuracy of A and only annotate the contents of B.) The chit-chat dialogue and reasoning dialogue are relatively unconstrained and freely conducted, without incorporating any specific human-machine interaction settings (Section 3.2). They necessitate only their memory and comprehension capabilities regarding contextual information, thereby minimizing the need for manual intervention. Consequently, for the responses of A in knowledge-grounded and task-oriented scenarios, where the contents do not conform to human language, we manually modify and have LLMs re-generate. Eventually, we obtain the complete dataset of multi-turn dialogue.
@@
We provide one output sample and more generation details in Appendix A.5. The statistical information of the whole benchmark is in Table 2.
@@
Avg. Response Length (Words) 13.2899